Steve's family traveled 3/4 the distance that Mike's family traveled. If Steve's family traveled 18 mi, how far did Mike's family travel?
Answer by checkley71(8403)

18=3/4 X OR 18 X(4/3)=X OR 72/3=X OR 24=X.
PROOF 3/4 OR .75 X 24 = 18.